OpenStudy (anonymous):

WILL GIVE MEDAL. BETTER LOOK OF EQUATION IN COMMENTS SIMPLIFY sqrroot10a times sqrroot30a

OpenStudy (anonymous):

\[\sqrt{10a}\times \sqrt{30a}\]

OpenStudy (anonymous):

PLEASE HELP. WILL GIVE MEDAL.

OpenStudy (anonymous):

\[\sqrt{300}\]

OpenStudy (anonymous):

and it is imperfect

OpenStudy (anonymous):

it was actually \[10\sqrt{3a}\]

OpenStudy (anonymous):

ok i was just taking a longshot at it

OpenStudy (anonymous):

tho, \[\sqrt{300}\] was one of the steps along the way lol

OpenStudy (anonymous):

ok lol